# Quasinormal ringing and Unruh-Verlinde temperature of the Frolov Black Hole

### v2

@@ VI Unruh Temperature
In this section we examine the Unruh temperature associated with the spacetime under consideration. This temperature reflects the surface gravity or, equivalently, the proper acceleration felt by an observer stationed at a fixed radial distance from the black hole. Such a property has also been explored in various alternative theories of gravity. The Unruh temperature can be expressed in terms of the red-shifted surface gravity, or more generally through covariant quantities involving the gravitational potential and the timelike Killing vector field. Previously, the Unruh temperature has been investigated in various setups Verlinde:2010hp ; Konoplya:2010ak
@@
Here $e^{\phi}$ represents the red-shift factor, which approaches unity at spatial infinity because $\phi=0$ as $r\to\infty$, provided the spacetime is asymptotically flat. The background geometry is assumed to be a static solution that admits a global timelike Killing vector $\xi^{\alpha}$.
@@
$a^{\alpha}=-g^{\alpha\beta}\nabla_{\beta}\phi.$ (34)
@@
Here $n^{\alpha}$ is a unit vector that is everywhere orthogonal to the timelike Killing vector $\xi^{\alpha}$. This equation can be rewritten using eq.(33) as Unruh:1976db
$T_{\text{Unruh}}=\frac{\hbar}{2\pi}\,e^{\phi}\sqrt{g^{\alpha\beta}\,\partial_{\alpha}\phi\,\partial_{\beta}\phi}.$ (36)
@@ Acknowledgments
The author is thankful to R.A Konoplya for useful discussions and for suggesting improvements on the draft. The author also expresses his gratitude to Pankaj S. Joshi and Parth Bambhaniya for giving the opportunity to visit ICSC, Ahmedabad University.

### v3

@@ VI Unruh Temperature
In this section, we examine the Unruh temperature associated with the spacetime under consideration. This temperature reflects the surface gravity or, equivalently, the proper acceleration felt by an observer stationed at a fixed radial distance from the black hole. Such a property has also been explored in various alternative theories of gravity. The Unruh temperature can be expressed in terms of the red-shifted surface gravity, or more generally through covariant quantities involving the gravitational potential and the timelike Killing vector field. Previously, the Unruh temperature has been investigated in various setups Verlinde:2010hp ; Konoplya:2010ak
@@
The quantity
$e^{\phi}=\sqrt{-\,g_{\alpha\beta}\,\xi^{\alpha}\xi^{\beta}}$ (37)
is the redshift factor associated with the static Killing field. It measures how the proper time of a stationary observer at radius $r$ relates to the coordinate time measured by an observer at spatial infinity. In an asymptotically flat spacetime, this factor approaches unity as $r\to\infty$, ensuring the standard normalization of the Killing vector.
The background geometry is assumed to be a static solution that admits a global timelike Killing vector $\xi^{\alpha}$. For the static and spherically symmetric spacetime considered here, we choose the timelike Killing vector associated with time translations, $\xi^{\mu}=(\partial_{t})^{\mu}$. Its covariant components are given by $\xi_{\mu}=g_{\mu\nu}\xi^{\nu}=(-f(r),0,0,0)$, where $f(r)$ is the metric function. This Killing vector is normalized such that $\xi^{\mu}\xi_{\mu}\to-1$ at spatial infinity.
@@
$a^{\alpha}=-g^{\alpha\beta}\nabla_{\beta}\phi,$ (38)
where
$\phi(r)=\frac{1}{2}\ln f(r)$ (39)
@@
Here $n^{\alpha}$ is a unit vector that is everywhere orthogonal to the timelike Killing vector $\xi^{\alpha}$. This equation can be rewritten as Unruh:1976db
$T_{\text{Unruh}}=\frac{\hbar}{2\pi}\,e^{\phi}\sqrt{g^{\alpha\beta}\,\partial_{\alpha}\phi\,\partial_{\beta}\phi}$ (41)
@@ Acknowledgments
The author would like to thank R. A. Konoplya for stimulating discussions and the anonymous referees for their valuable comments, which significantly improved the manuscript. The author also expresses his gratitude to Pankaj S. Joshi and Parth Bambhaniya for giving the opportunity to visit ICSC, Ahmedabad University.